OpenStudy (anonymous):

Brad sold candy bars and cookies for a fundraiser at school. Candy bars sold for $3 and cookies sold for $5. He sold a total of 20 items and made $76. How many candy bars did Brad sell?

OpenStudy (karatechopper):

20 items? thts a bit less for this question r u sure?

OpenStudy (anonymous):

yes im sure:)

OpenStudy (karatechopper):

ok...

OpenStudy (anonymous):

so do you know how to doit

OpenStudy (karatechopper):

not necesssairily...

OpenStudy (anonymous):

think I got it

OpenStudy (anonymous):

:D 12 candy bars were sold my friend.

OpenStudy (anonymous):

SO. check this out. you with me?

OpenStudy (anonymous):

3x+5y = 76$ right. and x+y = 20 right so solving for x in the second equation gets you x = 20-y put this in the top equation 3(20-y) +5y = 76 60-3y+5y = 76 2y = 16 y = 8 3x+5(8) = 76 3x+40 = 76 3x = 36 x = 12 y = 8

OpenStudy (anonymous):

yes that seems to make sense, thanks:) its basic substitution